The sum of two numbers is 29. Their difference is 7. What are the two numbers?
Answer by uma(370)

Let the two numbers be x and y.
Then x+y = 29 and x - y = 7
Adding the above two equations we get,
2x = 36
==> x = 36/2
==> x = 18
Substituting in the first equation,
18 + y = 29
==> y = 29-18
==> y = 11
Thus the two numbers are 18 and 11.
Good Luck!!!